The airplane, traveling at a constant speed of 60 m/s. is executing a horizontal turn. If the plane is banked at θ=20°, when the pilot experiences only a normal force on the seat of the plane, determine the radius of curvature ρ of the turn. Also, what is the normal force of the seat on the m pilot if he has a mass of 75 kg 3)
Answer #1

From the given question,

when pilot experiences only normal force, then

mgcos\theta=mv2/\rho*sin\theta

\rho=v2/(gtan\theta)

=(60*60)/(9.8*tan20)

=1009 m

Normal reaction= mgsin\theta+mv2/\rho*cos\theta

=75*9.8*sin20 + 75*60*60/(1009*cos20)

=536 N

Normal force of the seat is 536 N